OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

what does parallel mean?

OpenStudy (anonymous):

first off

OpenStudy (anonymous):

that there ryte by eachother

OpenStudy (anonymous):

hmm not exactly

OpenStudy (anonymous):

well i noe what it looks like iono how to put it in words

OpenStudy (anonymous):

they are any lines together however far apart but they will never intersect each other

OpenStudy (anonymous):

yea okay

OpenStudy (anonymous):

their distances must be the same all the way through and never insect

OpenStudy (anonymous):

yes inoe :)

OpenStudy (anonymous):

because x^2 will not intersect but it is not parallel but will not intersect

OpenStudy (anonymous):

anyhow

OpenStudy (anonymous):

okay:) lol

OpenStudy (anonymous):

you have to match 2 lines so one must have the same distance from the other

OpenStudy (anonymous):

yep got it

OpenStudy (anonymous):

first of off is it a positive slope or negative

OpenStudy (anonymous):

postive i think?

OpenStudy (anonymous):

well of course because original equation is y = 7x + 2 and you have to choose one that is parallel to it

OpenStudy (anonymous):

sooo wait what?!?!?!?

OpenStudy (anonymous):

i got confused

OpenStudy (anonymous):

if you see a positive number for the slope the slope is going to be positive

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that takes out D B those are negative slopes and they will intersect correct?

OpenStudy (anonymous):

yea because their negative slopes

OpenStudy (anonymous):

do you know how a positive slop and negative slope will intersect

OpenStudy (anonymous):

or what it looks like

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

k good

OpenStudy (candyme):

Start with y=mx+b and filll in what you know. y=7x+b Plug in your point to find b -3=7(1)+b -3=7+b -10=b Stick that value in. y=7-10

OpenStudy (candyme):

Sorry typo. should be y=7x--10

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

or you could do that :P

OpenStudy (anonymous):

aha thanks so much agin :D

OpenStudy (anonymous):

but in any case though

OpenStudy (anonymous):

hmmm?

OpenStudy (anonymous):

if you have a problem like this

OpenStudy (anonymous):

first notice the slope the must have the same slope because the are parallel

OpenStudy (anonymous):

if they do not they will intersect

OpenStudy (anonymous):

you can try it youself

OpenStudy (anonymous):

but ya

OpenStudy (anonymous):

okay aha thnks again :)